### v1

# Y-AR: A Mixed Reality CAD Tool for 3D Wire Bending

Abstract.
Wire bending is a technique used in manufacturing to mass-produce . CWire bending machines like the DIWire by Pensalabs111https://www.pensalabs.com/diwire-pro have made this process accessible for personal fabrication. However, such machines are controlled using Computer Aided Manufacturing (CAM) software which is hard to use, making custom design challenging. We present Y-AR, a Computer Aided Design (CAD) interface for 3D wire bending. Y-AR uses mixed reality to let designers create structures that physically connect to objects in the environment (such as the mount shown in Figure 1c). The interface incorporates springs as design primitives which (1) apply forces to hold objects, and (2) counter-act dimensional inaccuracies inherently caused by mid air modeling and measurement errors in AR. We demonstrate design workflows to design and fabricate a range of mechanisms designed in Y-AR as well as structures made using free-hand design tools. In our usability evaluation, all 12 participants successfully designed and fabricated a functional bottle holder with Y-AR. 10 out of 12 participants felt that the system aided their design process, rating it above 7 out of 10.
1. Introduction
Wire bending is a common industrial mass-manufacturing process used to create springs, clips, mounts, and braces. In sculpting, it is used to provide internal support for incomplete structures, particularly in the form of armatures. Most wire bending occurs in mass-production with high-volume, low-mix production. Today, there is a growing trend towards custom wire bending, where each object is heterogeneous and crafted individually rather than the mass manufacture of wire objects with uniform geometry.This custom process is predominantly manual. In the medical sector, it is used to create retainers and braces (Alassiry, 2019), and in puppetry, it enables the production of actuated characters (Xu et al., 2018), both are examples of low-volume, high-mix production (e.g., few unique products).
Machines like the DIWire by Pensalabs have made this process accessible for personal fabrication (Baudisch and Mueller, 2017). Hobbyists have embraced the technology to make connectors, holders, signs, marble tracks, and sculptures. While software exists to operate the machines (e.g., VGPNEXT222https://www.blmgroup.com/en-us/software/vgpnext, Bend-Tech333https://www.bend-tech.com/, and BendPro SIM444https://www.currenttech.com/tube-bending-software/bendpro-sim/), they naturally inherit design approaches from industrial manufacturing. Users apply sequential combinations of bend, rotation, and extrusion operations. Such low-level, machine-specific commands (common in Computer Aided Manufacturing (CAM)) make it hard for users without technical expertise to design structures for wire bending.
Researchers have built software converting existing 3D models to bend paths (Huang et al., 2023; Hsiao et al., 2018). These tools let hobbyists create intricate wire structures based off 3D models, however modifying models after conversion remains challenging, requiring round-trips of conversion. These design tools are agnostic to the strengths and weaknesses of wire bending. Often there are multiple ways to convert 3D models to paths for wire bending, WireRoom (Yang et al., 2021) eloquently addresses this through a user interface that offers multiple generated wire bending paths to choose from. However, to date, there are no Computer Aided Design (CAD) tools designed to support wire bending.
We present Y-AR, a CAD tool for 3D wire bending in mixed reality. It is a fabrication-aware-design (Bermano et al., 2017) tool that lets designers create 3D structures optimized for CNC wire bending. Y-AR overlays designed 3D structures on users’ physical environments in mixed reality to create objects fitting within the physical environment of the designer (examples shown in Figure 2). Y-AR uses design primitives such as mounts, and holders, to connect structures to physical artifacts. Y-AR leverages the elasticity of wires to create mechanisms that apply force to hold objects in place. This springiness compensates for the dimensional inaccuracies inherent to XR-based modeling and hand-tracking. It automatically generates machine instructions for wire benders and assembly instructions for users. We present the design space of primitives consisting of clamps and mounts to hold objects with shapes such as cylinders, boxes, spheres, and higher-order polyhedra. We furthermore demonstrate free-form design (where users draw structures in mid air) by modeling a custom marble track. In our usability evaluation, all 12 participants successfully designed and fabricated a functional bottle holder with Y-AR. 10 out of 12 participants felt that Y-AR aided their design process, rating it above 7 out of 10.
2. RELATED WORK
This work builds on research on wire bending, fabrication-aware design, and computer-aided design in mixed reality.
2.1. Wire Bending
Research on wire bending has focused on aesthetic and functional wire bending. For aesthetic approaches, WrapIt created jigs to speed up manual wire bending for jewelry based on 2D patterns (Iarussi et al., 2015). Fabricable 3D Wire Art similarly generate jigs however they base it on 3D models, and images enabling more complex designs (Tojo et al., 2024). With the advent of CNC wire bending machines, there has been an increased focus on advanced wire bending by converting 2D line drawings (Hsiao et al., 2018) or 3D models (Huang et al., 2023) to machine instructions for wire bending. Yeh et al. (2022) propose an algorithm to create 3D wire art with the minimum number of continuous wires. In terms of user interfaces to generate 3D wire art, WireRoom provides an interface to offer multiple options to generate wire bending paths (Yang et al., 2021) allowing users to select their preferred result. PM4WireArt offers users more freedom through a parametric modeling interface designed for conceptual wire art design (Li, 2024). However, instead of focusing on traditional wire bending—where a single wire is bent and rotated into shapes—PM4WireArt treats wire as rigid rods assembled into structures, resembling welded frameworks. This approach facilitates complex design creation highlighting a different use of wire in structural art.
Several projects explore functional instead of purely aesthetic properties of wire bending. For instance, Bend-forming extends the application of wire bending to aerospace engineering. Their work involves the generation of truss structures and explores the concept of deploying wire benders in space to construct large-scale, static structures (Bhundiya and Cordero, 2023). More down to earth, WireFab generates internal skeletons for existing 3D models, serving as static frameworks with some movement flexibility (Liu et al., 2017). Finally, Bend-it bends kinematic structures by taking key frames of a virtual structure designed by wires as input and generating kinematic mechanisms to animate between these frames, utilizing the springiness and dynamic properties of the wire (Xu et al., 2018). Similar to the aesthetic applications, these fabrication techniques are based on conversion of 3D structures and modification of pre-defined parametric models.
2.2. Fabrication-Aware Design
Fabrication-aware design is a branch of research in computer graphics that considers the constraints and possibilities of fabrication processes in design environments. This approach ensures that resulting artefacts are manufacturable and functional throughout the design process. Fabrication-aware design has implications in many fields of computational fabrication such as laser cutting, 3D printing, and other fabrication processes.
In the context of laser cutting,kyub is a 3D modeling environment that allows users to design models for laser-cutting in 3D while the software handles the conversion to 2D plates by placing joints for the specific material and machine at hand (Baudisch et al., 2019). FastForce safeguards the structural integrity of such laser cut models (Abdullah et al., 2021) and FoolproofJoint adjusts resulting joints to be optimal for assembly (Park et al., 2022). HingeCore leverages both laser cutting and specific material properties to create models that are aesthetically pleasing and easy to assemble (Abdullah et al., 2022). Fabricaide allows users to preview fabrication process constraints (such as issues in nesting or joint layout) during the design phase (Sethapakdi et al., 2021). FlatFab similarly supports design for fabrication for laser cutting by focusing on a specific assembly technique of intersecting planar sections (McCrae et al., 2014).
In the realm of 3D printing, fabrication-aware design tools have also made significant strides. Bridging the gap introduces a method to efficiently design stable supports using bridges and pillars (Dumas et al., 2014). Stress relief detects structural weaknesses and provides tools for users to adjust their model accordingly (Stava et al., 2012). To increase print efficiency, WirePrint offers a rapid prototyping method by generating wireframe structures from existing 3D models (Mueller et al., 2014), and PackMerger optimizes the segmentation and packing of printed parts to save material and print time (Vanek et al., 2014). Others have brought material specific properties to fabrication-aware design, Wang et al. (2015) reduce fabrication time while maintaining visual quality through adaptive slicing and segmentation techniques. Bickel et al. (2010) use data-driven approaches to create complex, heterogeneous materials with customized deformation properties . Schumacher et al. (2015) design micro-structures to achieve variable elasticity using a single stiff material, expanding supported design applications.
While significant advances have been made in developing fabrication-aware design tools for laser cutting and 3D printing, it is not limited to these processes. In textile design, Plushie allows non-professional users to create plush toys by generating 2D patterns from 3D models (Mori and Igarashi, 2007). For inflatable structures, Skouras et al. (2014) developed a tool for optimizing flat panel patterns for inflatable structures, ensuring accurate design through physics-based simulations. In the project Beyond developable, researchers explore using auxetic materials and conformal geometry to create doubly-curved surfaces from flat pieces (Konaković et al., 2016).
Wire bending remains under explored in terms of fabrication-aware design. Miguel et al. (2016) focus on creating structurally stable wire sculptures rather than serving as a design tool. Wire mesh design explores the broader design potential of wire based materials (Garg et al., 2014). Specialized tools for wire bending could unlock new design possibilities and elevate the complexity and functionality of wire-based structures.
2.3. Computer Aided Design in Mixed Reality
Extended Reality (XR) CAD technologies offer advantages for digital fabrication. Tools like pARam, allow in-situ previewing of designs (Stemasov et al., 2024), Mix&Match enables modeling-free personal fabrication with real-world dimensions (Stemasov et al., 2020), and MixFab, integrates virtual CAD environments and real-world objects to create a mixed-reality environment for personal fabrication, to lower the barrier to entry for broad adoption (Weichel et al., 2014). Additionally, commercial products such as Gravity Sketch555https://www.gravitysketch.com/ demonstrate the practical advantage of XR CAD in design, allowing for quick creation of 3D models, even if they lack the precision of traditional CAD tools. Microsoft’s Maquette666https://store.steampowered.com/app/967490/Microsoft_Maquette/, facilitates spatial prototyping by combining various tools such as surface snapping, hull brush, bounding manipulation, to manipulate 3D content.
XR presents challenges as well. One major issue is the difficulty in accurately obtaining environmental information, and users interacting in mid-air resulting in misalignment between the virtual and physical environments. Even though research has begun exploring ways to align virtual environments with the real world, such as SnapToReality, which proposes a semi-automatic alignment strategy for registering virtual and real spaces, their work relies on scene content and environmental constraints (e.g., edges and surfaces) to anchor virtual content (Nuernberger et al., 2016). Another approach, presented by Fotouhi et al. (2020) requires specific headsets and additional equipment.
Another challenge lies in tracking or gesture detection. Efforts to define hand gestures face limitations, such as high similarity among gestures leading to recognition errors and limited available information (Zhou et al., 2016). Some solutions, like high-quality tracking devices, exist but are often prohibitively expensive and not accessible for everyday home use.
Y-AR leverages the strengths of XR CAD, such as the ease of obtaining real-world dimensions, while mitigating its shortcomings. By combining XR CAD with wire as a material, we leverage the springy nature of materials to design mechanisms that are robust to design, measurement, and fabrication inaccuracies, a concept similar to SpringFit (Roumen et al., 2019) used in mechanical engineering. This approach allows us to create adaptable mechanisms in Y-AR that maintain stability and functionality despite misalignment, addressing limitations related to precise alignment.
3. Walk-through of Y-AR
Figure 3 shows a walk-through of a user making a mount for a scotch tape extruder on a pegboard. (a) The user puts on their Meta Quest 3 and invokes Y-AR. In Passthrough mode, the user selects the type of connectors. (b) Y-AR places a generic pegboard connector and scales this to the detected gesture, by measuring the distance between the fingers. (c) The user then selects another connector using interface shown in (a) and grabs the scotch tape extruder, to generate a mount which fits the size of the handle. (d) The user links the two connectors to one another to complete the design and fabricate (e) the resulting mechanism with the wire bender.
After design, users export their structure. Export generates a CSV file for the wire bender (shown in Figure 4 as well as assembly instructions for the user in VR. Y-AR aims to make structures out of a single wire, however often this is not feasible, upon which Y-AR exports multiple wires to bend and instructions in VR which guide users through the process of assembling the wires into a 3D structure, as shown in Figure 4b. To rigidly connect wire segments, we use a Crimping Tool Kit777https://www.amazon.com/dp/B0C7V6K4X3 (other similar kits will do too). This allows for secure and precise joining of wires by crimping connectors onto the ends of the wire segments.
4. contributions, benefits, and limitations
We make two core contributions: (1) guidelines for fabrication-aware wire bending and (2) an approach for leveraging material springiness to overcome accuracy issues inherent to mid-air interaction and tracking in XR. We present both contributions as separate sections in the paper as well as integrated in the specific implementation of a first CAD tool for wire bending, we call Y-AR.
(1) We systematically explore the design space of connectors made with wire bending, through this we identify technical guidelines for functional wire bending.
(2) The insight to leverage designed springiness to overcome modeling and tracking inaccuracies serves as a solid basis to implement fabrication-aware CAD tools. We demonstrate this at the example of wire bending in Y-AR. To validate the design of Y-AR, we conduct a usability study.
Limitations include the focus on one specific material used for wire bending which is 1.6 mm red nylon coat wire, while the ideas and methods translate to other materials, it would require re-configuration of some tools in Y-AR. Furthermore, while Y-AR avoids cutting wires (resulting in manual assembly), this algorithm is far from optimal. To the best of our knowledge, most wire benders take csv files containing a list of coordinates as input, however, high-end industrial machines may use more advanced proprietary formats, currently not supported by Y-AR. Y-AR currently supports a series of useful modeling tools, however we realize this set is far from exhaustive. We open-source the code base and make it easy to extend to such formats or other wire types.
5. Design of Y-AR
Y-AR follows a verb-noun interaction model. The Y-AR interface consists of a top-level menu with different actions: creating connectors, mid-air drawing, and creating marble tracks. Next to that, it comes with basic editing tools such as linking structures, deleting, and modifying or moving components and a generic export function. As illustrated by Figure 2, Y-AR allows users to create models that are hard to bend by hand and challenging to produce with CAM tools.
5.1. Create connector tool
The create connector tool lets designers create wire-bending structures to connect to existing objects or to features in the context of the user. The workflow for this tool is as follows:
Users select a connector type and press the gesture button to enable gesture detection (this is not ”always on” to avoid undesired detection of gestures). They simply grab the object the way they want to connect, Y-AR detects the grabbing gesture and extracts dimensions for the required connector. Dimensions include the distance between the index finger and thumb to control the size of the generated connector, the hand’s orientation, and the palm’s position to control the placement of the generated connector, find specific details in section 7.1.
The generated connector shows up in the XR model as a preview of what the wire structure will look like. Upon placement, the modify tool is triggered so users can immediately move, rotate and scale the connector as needed.
As mentioned before, Y-AR requires users to choose the connector type. This allows it to identify the right parameters for each type of connector. Below we specify how these connector types relate to their respective gestures.
The table edge connector measures the table edge’s thickness using the vertical distance between the thumb tip and palm (see Figure 5 b).
The cylindrical connector determines the cylinder diameter based on the distance between the thumb and index finger during a grabbing gesture (see Figure 5 c).
And the pegboard connector sets the distance between connector legs using the spread of the index and middle fingers (see Figure 5 a).
5.2. Mid-air drawing tool
The mid-air drawing tool allows users to freely draw lines in mid-air, generating paths with more flexibility. As shown in Figure 6, users can draw in various contexts, such as (a) mid-air, (b) on the human body, (c) on the surfaces of existing objects like lampshades for decorative purposes, or (d) create mesh structures based on dress forms. These capabilities enable the creation of unique wire structures, such as wire-wrapped jewelry, which benefit from lightweight connections, high flexibility, and aesthetic customization.
In our implementation of Y-AR, users use the Logitech MX Ink, pen-based tool for Mixed Reality (MR) with high-accuracy drawing functions, to start drawing a line the user press the button in the MX Ink and draws in mid-air, each line end upon releasing the button. Alternatively, users perform a pinch gesture to start drawing a line until the pinch gesture is ended without the need of the MX Ink. Users freely draw lines to model their desired structures in 3D.
5.3. Marble track tool
To create more deliberate functional structures, we let users design structures using what we call Shape Modules. To exemplify, we implemented a Marble Track tool demonstrated in Figure 7. Users create freehand structures consisting of pre-defined modules. This tool is developed based on the Mid-Air Drawing Tool, and its basic workflow is similar.
When users sketch a line for a marble track, the system automatically generates not just a single wire, but four parallel lines to form a robust track. It occasionally injects support towers consisting of vertical truss structures as well as small 3D printed components to hold the wires together. Additionally, based on the track’s length, Y-AR places additional support structures at midpoints along the track to prevent sagging and maintain structural integrity under the weight of the marbles. Our current marble track modules are optimized for standard 1.6 cm in diameter marbles, it would be relatively easy to extend to other sizes and types.
5.4. Other editing tools
Additionally we provide a set of generic editing tools to modify designs. These tools apply to all structures and content in the scene. They consist of a delete tool to delete segments or lines; a modify tool, to move scale or rotate segments or lines, and a link tool to link individual line segments or connectors to one another. Finally, there is one export function to convert the designed wires to machine instructions (csv) and assembly instructions (AR preview).
5.5. Gesture training
Traditionally, gesture-based interfaces are not the most discoverable (Zhou et al., 2016). However, to create connectors, Y-AR leverages gesture interaction by having users demonstrate how the structure should mount objects. In doing so, we build on work by Pei et al. (2022), which lets users’ hands become virtual objects by imitating the objects themselves. To aid users in adopting these gestures, every tool in Y-AR includes an interactive on-boarding mode, as shown in Figure 8.
6. implementation
Y-AR is a Unity application, it implements 3D and UI interactions via Unity input events and includes customizable 3D UI elements for on-boarding, spatial data collection, and Passthrough mode setup. The chosen tech stack, including Unity version 2022.3.8f1, XR Interaction Toolkit (version 2.5.1), XR Hands (version 1.4.1), XR Core Utilities (version 2.2.3), XRC Line Tool (version 0.1.0), and XR Core (version 0.1.4), was selected for its comprehensive support for mixed reality development and ease of integration with various XR features.
Y-AR achieves hand gesture tracking by using the Unity XR Hands package. Gesture recognition translates hand movements into CAD system actions, with real-time feedback provided by the XRC Line Tool from the XRC core, which is used for creating, rendering, and manipulating lines in the project.
Y-AR achieves the transformation of spatial data into a wire-bending path through a dynamic tracing system that accurately places and scales connectors. This system measures distances between specific hand joints to proportionally adjust the sizes of virtual objects, ensuring precise and adaptable designs. Y-AR achieves mid-air drawing by using the controller’s position as input, visualizing this with the XRC Line Tool. The system utilizes compatible controllers to track spatial movements, which are then translated into wire-bending paths and exported as drawing data.
The bending paths designed in Y-AR are exported as CSV files, containing a list of coordinates along the path. This format is chosen for its simplicity and compatibility, allowing the files to be easily read and interpreted by wire bending software. This ensures that the designs can be accurately translated into physical forms by the wire bender.
7. Fabrication-Aware Design for Wire Bending
In this section, we systematically explore fabrication-aware design for wire bending through the design of different types of connectors, and the limitations of materials and fabrication machines.
7.1. Connector Design
We designed a series of connectors using wire bending to connect different types of objects. As shown in Figure 9, the target objects are defined as collective variances of geometric primitives (including: (a) a hanging rectangular shape, (b) a cylindrical shape supported on the bottom, (c) a hanging cylindrical shape, (d) a horizontally oriented rectangle, and (e) spherical shapes).
Based on these designs, we developed several design guidelines to guarantee successful wire bending with a numerical bending machine: (1) avoiding planar bends, (2) minimum bending angle to overcome plastic deformation, (3) constraining 3DoF, (4) collision detection.
7.2. Avoid planar bends
In our design tests, we often find that when performing multiple bends (e.g., three consecutive 90-degree bends), the wire may collide with the machine or the extruded wire itself, leading to wire bending failure. To address this issue, our strategy is to introduce rotations at the nodes, particularly when designing planar structures as shown in Figure 10 a. By incorporating rotations, as illustrated in Figure 10 b, we ensure that previously bent sections of the wire and subsequent bends are not on the same plane, thereby avoiding collisions and ensuring successful wire bending.
7.3. Minimum bending angle to overcome plastic deformation
To plastically deform wires (bend them permanently), it is crucial to overcome the elastic regime of the material. Small bends simply bend the wire and the wire bends back. To overcome this, we identify a minimum required bending angle and over-extend bends where needed.
In our design process, we tested bending angles from 5, 15, 25, 35, 45… up to 85 degrees (within a 90-degree range) and compared these with the actual bending results. This approach helped us effectively avoid the issue of ensuring that small bending angles in the design would still result in a noticeable bend.
As shown in the Figure 11, when the machine was set to bend at 15 degrees, the actual bending angle was significantly reduced. This finding indicates that in our designs, we should avoid specifying angles smaller than or around 15 degrees.
7.4. Constraining 3DOF
In our connector design, we primarily focus on constraining the three degrees of freedom (minus rotation) of objects. This means that to fix an object in a spatial position, all planar degrees of freedom need to be restricted. We use three main strategies to achieve this:
Gravity-Based Holders. These connectors rely on gravity to hold objects in place. As shown in Figure 9 c, cylindrical objects, such as water bottles, hang on the holder using their weight. The drawback of this method is that if the holder’s orientation changes, the fixed object may no longer be secured.
Clamp-Type Holders. These connectors use force in one direction to secure the object, while the object’s inherent limitations prevent movement in another direction. Examples include Figures 9 a and d. The advantage of this method is that, ideally, if the force is strong enough, the object remains fixed even if the connector’s orientation changes. However, as seen in Figure 12 a, if the force (both positive and negative Y direction) is insufficient, the object might move along the X direction perpendicular to the applied force.
Combination Holders. These connectors combine the principles of both gravity-based and clamp-type holders. An example is Figure 9 b (cafe cup holder) and Figure 9 e. The cafe cup holder clamps the coffee cup to prevent it from spilling, applying force in both positive and negative Y-directions (As shown in 12 b). Gravity effectively restricts movement in the vertical direction. This combined approach ensures stability and minimizes the risk of the object becoming unsecured due to changes in orientation.
7.5. Collision detection
Users lack knowledge about the machine’s limitations to bend wires. Issues can arise when lines are too close to each other or when the angle between them is too small, potentially causing machine collisions during bending.
To address this, we implemented basic collision detection. Our algorithm sets a minimum threshold angle that the wire bender can fabricate. Since all lines are drawn by connecting multiple points, the algorithm separates all points and combines every pair of points into a line segment. By calculating the angles between two such segments lines, the algorithm determines whether their angle exceeds the threshold. If it is, those segments are flagged as conflicts.
8. USER STUDY
To validate our implementation of Y-AR we conduct a basic usability study, demonstrating that non-technical users can indeed reliably design and fabricate mounting structures with Y-AR.
8.1. Participants
We recruited 12 participants from our university (number of female students = 7, age mean = 24.33, SD = 1.83). Most participants (10) reported having beginner proficiency with CAD tools, and only a few (3) indicated they engaged in wire bending activities before.
8.2. Hardware
We used the Meta Quest 3 as the study apparatus. We fabricated the designed structured during the study with the C64 CNC Wire Bender from PensaLabs. As input controller, we used the MX INK from Logitech.
8.3. Task
Each participant performed two tasks with Y-AR. Task 1 involves designingsm to connect a bottle to a table edge (see Figure 14). Task 2 is an open-ended exploration of Y-AR.
8.4. Metrics
To assess the usability, we selected specific questions from standardized tests: the System Usability Scale (SUS) and the User Experience Questionnaire (UEQ). Additionally, the mechanical performance of the resulting structures was evaluated by mounting a bottle to determine whether it could be securely supported.
8.5. Procedure
This study lasted 60 minutes in a controlled lab setting. Participants performed two tasks and were given 5 minutes to familiarize themselves with the software through a training video. In the first 15-minute task, they conducted task 1. If participants did not complete their modeling within the allocated time, their modeling was marked as incomplete. If they finished early, they could inform the researcher and proceed to the next stage. They fabricate the structure with help from the researchers and assemble their designs. They then proceeded with task 2 in another 15-minute block.
Finally, participants filled out our usability survey and provided feedback on their experiences in a semi-structured interview. All participants completed the study within 60 minutes.
8.6. Results
User Experience. All ratings in this evaluation are based on a 10-point scale and result is shown in the Figure 15. Out of the 12 participants, 9 indicated they would like to use Y-AR frequently, rating it more than 7. Ten participants felt confident using Y-AR, also rating it more than 7, and 11 believed most people would quickly learn to use Y-AR, with ratings above 7. Additionally, 10 participants enjoyed using hand gesture modeling, rating this aspect above 7, and 10 enjoyed the design process overall, with ratings above 7; 2 participants rated it at 6. The AR environment was found useful for obtaining real-world dimensions by all participants. Eight participants appreciated the gestural icons, rating them more than 7. Only three participants found Y-AR complex, rating this aspect above 3, and only two participants indicated they needed to learn many things before starting, rating this aspect above 7. Lastly, 11 participants considered the environment of the room in their design, with ratings above 7.
Technical Performance. As shown in Figure 16, the accuracy of hand gesture recognition was rated positively by 6 participants, with scores above 7, and 3 participants rated it at 6. The mechanical connector design met the expectations of 10 participants, who rated it above 7, while 2 participants rated it at 6. Despite only 4 participants having a clear strategy in mind when designing the connector, all were able to design a connector successfully using Y-AR, with 4 rating their strategy above 7 and 3 at 6. Eleven participants reported a good sense of size, with ratings above 7, and similarly, 11 participants felt confident in the spatial positioning while using Y-AR. Furthermore, all participants believed that the connectors, once manufactured, would look as expected, rating this aspect above 7, and 11 participants felt the connectors would function as intended, also rating above 7. Additionally, all participants acknowledged that Y-AR helped them design a functional object.
Areas for Improvement. As shown in Figure 17, the results also highlighted some areas for improvement. Three participants felt they would need technical support to use Y-AR effectively, rating this need less than 3, and 2 participants rated it at 4, suggesting that Y-AR may require a more intuitive on-boarding process. Additionally, four participants reported experiencing arm fatigue, with scores indicating no arm fatigue below 4.
ubsectionDiscussion
Ease of Use and Intuitiveness.
Participants found Y-AR straightforward and easy to use, highlighting its simplicity and intuitive design. For example, P12 mentioned, ”It was easy to make… So it’s really useful for me to use pre-made wire same.” P11 appreciated Y-AR’s straightforwardness: ”I think using the gesture could easily reflect the actual size and is very straightforward without the hassles of using measuring rulers to measure the wires and then put them into the CAD model.”
P10 compared Y-AR favorably to other drawing tools, stating, ”Unlike some other drawing tools, where I have to select different pens, thickness, and colors, I can just draw and get what I want right away.” P8 initially felt overwhelmed due to unfamiliarity with the XR environment but noted, ”After some kind of introduction, I was getting familiar with that. I was pretty comfortable during the experiment.”
P7 found the freehand drawing feature intriguing, saying it was ”very interesting to be able to draw freely in the space around me.” However, P6 highlighted a potential issue, expressing that while Y-AR is simple, ”from my perspective, it doesn’t seem that user-friendly… especially with the headset’s operation logic and gesture recognition,” suggesting a need for better support in understanding XR operation systems.
P3 enjoyed Y-AR’s features, remarking, ”I really enjoy the function of the tool as you can draw anything you want in the VR world, which is pretty interesting, amazing.” They also found it helpful for connecting different parts and having dimensions available, which they described as ”simple and helpful.” Finally, P1 succinctly captured Y-AR’s appeal, stating, ”The tool is very intuitive, allowing users to quickly understand what they should do.”
Real-Time Feedback and Visualization.
Participants appreciated Y-AR’s ability to provide immediate feedback on designs and see changes in the structure in real-time, finding this feature valuable for making quick adjustments and ensuring their designs met expectations. For instance, P6 noted, ”This feature is particularly helpful for designing objects that have practical use in real life. It provides a significant visual aid during the design process.”
P6 highlighted the benefits of using XR devices for design, contrasting it with traditional tools like Blender: ”Using tools like Blender, I found it difficult to visualize 3D space. However, with VR, seeing a 3D image projected into the space provides significant help, especially in practical design applications. For instance, even if I know the dimensions of a table, seeing it visually in the space makes a big difference.”
Potential for Practical Application.
Participants saw significant potential for Y-AR in various practical applications. They envisioned using it for creating customized holders or organizers, particularly in home and kitchen settings. P11 noted, ”I think for wires like charging cable… we definitely need organizers to make everything in order, so I think that is a good use case scenario.” The tool was also recognized for its potential in artistic and creative applications. P9 discussed this, mentioning, ”I can definitely imagine this to be for not their practical but more artistic purpose… just hanging on your wall simply.”
Participants highlighted several applications, such as phone holders and organizers for items. P8 shared, ”For normal people, they would like to use a phone holder. It’s a very common use case. And also, something like attaching a bottle to the table or similar applications.” P8 also suggested a feature enhancement: ”Probably we can add something like if you have a phone and it’s set at a certain angle, it could automatically generate a structure underneath to support it.”
Feature Requests
Participants provided several suggestions for enhancements and new features in Y-AR to improve the user experience and functionality in the future. One key recommendation was the inclusion of more templates, which would assist users in creating more complex or functional designs without needing detailed technical knowledge. As one participant expressed, ”Having more predefined structures would be really helpful” (P12). Another participant highlighted the potential benefit of a ”measurement feature” similar to those in CAD tools, which would allow users to easily see dimensions and scale objects (P7).
9. CONCLUSION AND FUTURE WORK
We presented Y-AR, a CAD tool for 3D wire bending. This fabrication-aware environment allows designers to create structures with good wire bending in mind. We see our guidelines for design for wire bending as contributions independent of the specific implementation of the Y-AR environment. The core insight to leverage the springiness of the material to overcome modeling inaccuracies allows Y-AR to rely on much less precise input techniques such as mid-air gestures in XR. This XR environment, in turn, contributes to better design by guiding users to make objects that fit in with their context of use, we imagine this central idea will allow developers to build other fabrication-aware CAD tools for other fabrication techniques than wire bending. Our user evaluation indicates that Y-AR indeed allows users to create functional structures with wire bending that would have been hard to accomplish with existing CAM tools for wire bending.
Y-AR will be available as an open source tool for other researchers to extend. It would be interesting to extend Y-AR with new types of tools, or consider facilitating design of structures that are complemented by other fabrication techniques than wire bending alone. Related work has shown that wire bending is well equipped to provide structure for other fabrication processes. We are also interested in running more longer-term deployment to see how users benefit from Y-AR over an extended amount of time. And as stated before, the framework of Y-AR combining XR and springiness could form a backbone for other future fabrication-aware design environments.
Acknowledgements.
We thank the XR Collaboratory at Cornell Tech for their generous sponsorship of this research in the form of the inaugural XR Prototyping Grant.References
- (1)
- Abdullah et al. (2022) Muhammad Abdullah, Romeo Sommerfeld, Bjarne Sievers, Leonard Geier, Jonas Noack, Marcus Ding, Christoph Thieme, Laurenz Seidel, Lukas Fritzsche, Erik Langenhan, Oliver Adameck, Moritz Dzingel, Thomas Kern, Martin Taraz, Conrad Lempert, Shohei Katakura, Hany Mohsen Elhassany, Thijs Roumen, and Patrick Baudisch. 2022. HingeCore: Laser-Cut Foamcore for Fast Assembly. In Proceedings of the 35th Annual ACM Symposium on User Interface Software and Technology. ACM, Bend OR USA, 1–13. https://doi.org/10.1145/3526113.3545618
- Abdullah et al. (2021) Muhammad Abdullah, Martin Taraz, Yannis Kommana, Shohei Katakura, Robert Kovacs, Jotaro Shigeyama, Thijs Roumen, and Patrick Baudisch. 2021. FastForce: Real-Time Reinforcement of Laser-Cut Structures. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems. ACM, Yokohama Japan, 1–12. https://doi.org/10.1145/3411764.3445466
- Alassiry (2019) Ahmed M Alassiry. 2019. Orthodontic retainers: A contemporary overview. J Contemp Dent Pract 20, 7 (2019), 857–862.
- Baudisch and Mueller (2017) Patrick Baudisch and Stefanie Mueller. 2017. Personal fabrication. Now Publishers.
- Baudisch et al. (2019) Patrick Baudisch, Arthur Silber, Yannis Kommana, Milan Gruner, Ludwig Wall, Kevin Reuss, Lukas Heilman, Robert Kovacs, Daniel Rechlitz, and Thijs Roumen. 2019. Kyub: A 3D Editor for Modeling Sturdy Laser-Cut Objects. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems. ACM, Glasgow Scotland Uk, 1–12. https://doi.org/10.1145/3290605.3300796
- Bermano et al. (2017) Amit H. Bermano, Thomas Funkhouser, and Szymon Rusinkiewicz. 2017. State of the Art in Methods and Representations for Fabrication-Aware Design. In Eurographics State of the Art Reports.
- Bhundiya and Cordero (2023) Harsh G Bhundiya and Zachary C Cordero. 2023. Bend-forming: A CNC deformation process for fabricating 3D wireframe structures. Additive Manufacturing Letters 6 (2023), 100146.
- Bickel et al. (2010) Bernd Bickel, Moritz Bächer, Miguel A. Otaduy, Hyunho Richard Lee, Hanspeter Pfister, Markus Gross, and Wojciech Matusik. 2010. Design and fabrication of materials with desired deformation behavior. In ACM SIGGRAPH 2010 Papers (Los Angeles, California) (SIGGRAPH ’10). Association for Computing Machinery, New York, NY, USA, Article 63, 10 pages. https://doi.org/10.1145/1833349.1778800
- Dumas et al. (2014) Jérémie Dumas, Jean Hergel, and Sylvain Lefebvre. 2014. Bridging the gap: automated steady scaffoldings for 3D printing. ACM Trans. Graph. 33, 4, Article 98 (jul 2014), 10 pages. https://doi.org/10.1145/2601097.2601153
- Fotouhi et al. (2020) Javad Fotouhi, Tianyu Song, Arian Mehrfard, Giacomo Taylor, Qiaochu Wang, Fengfan Xian, Alejandro Martin-Gomez, Bernhard Fuerst, Mehran Armand, Mathias Unberath, et al. 2020. Reflective-ar display: An interaction methodology for virtual-to-real alignment in medical robotics. IEEE Robotics and Automation Letters 5, 2 (2020), 2722–2729.
- Garg et al. (2014) Akash Garg, Andrew O. Sageman-Furnas, Bailin Deng, Yonghao Yue, Eitan Grinspun, Mark Pauly, and Max Wardetzky. 2014. Wire mesh design. ACM Trans. Graph. 33, 4, Article 66 (jul 2014), 12 pages. https://doi.org/10.1145/2601097.2601106
- Hsiao et al. (2018) Kai-Wen Hsiao, Jia-Bin Huang, and Hung-Kuo Chu. 2018. Multi-view wire art. ACM Trans. Graph. 37, 6, Article 242 (dec 2018), 11 pages. https://doi.org/10.1145/3272127.3275070
- Huang et al. (2023) Huiguang Huang, Dong-Yi Wu, Hanh Le, Po-Chih Chen, Shih-Syun Lin, and Tong-Yee Lee. 2023. Learning to Generate Wire Sculpture Art from 3D Models. In SIGGRAPH Asia 2023 Posters (Sydney, NSW, Australia) (SA ’23). Association for Computing Machinery, New York, NY, USA, Article 25, 2 pages. https://doi.org/10.1145/3610542.3626124
- Iarussi et al. (2015) Emmanuel Iarussi, Wilmot Li, and Adrien Bousseau. 2015. WrapIt: computer-assisted crafting of wire wrapped jewelry. ACM Trans. Graph. 34, 6, Article 221 (nov 2015), 8 pages. https://doi.org/10.1145/2816795.2818118
- Konaković et al. (2016) Mina Konaković, Keenan Crane, Bailin Deng, Sofien Bouaziz, Daniel Piker, and Mark Pauly. 2016. Beyond developable: computational design and fabrication with auxetic materials. ACM Trans. Graph. 35, 4, Article 89 (jul 2016), 11 pages. https://doi.org/10.1145/2897824.2925944
- Li (2024) Wanwan Li. 2024. PM4WireArt: A Scriptable Parametric Modeling Interface for Conceptual Wire Art Design Using PM4VR. In 2024 7th International Conference on Information and Computer Technologies (ICICT). 338–343. https://doi.org/10.1109/ICICT62343.2024.00061
- Liu et al. (2017) Min Liu, Yunbo Zhang, Jing Bai, Yuanzhi Cao, Jeffrey M. Alperovich, and Karthik Ramani. 2017. WireFab: Mix-Dimensional Modeling and Fabrication for 3D Mesh Models. In Proceedings of the 2017 CHI Conference on Human Factors in Computing Systems (Denver, Colorado, USA) (CHI ’17). Association for Computing Machinery, New York, NY, USA, 965–976. https://doi.org/10.1145/3025453.3025619
- McCrae et al. (2014) James McCrae, Nobuyuki Umetani, and Karan Singh. 2014. FlatFitFab: interactive modeling with planar sections. In Proceedings of the 27th Annual ACM Symposium on User Interface Software and Technology (Honolulu, Hawaii, USA) (UIST ’14). Association for Computing Machinery, New York, NY, USA, 13–22. https://doi.org/10.1145/2642918.2647388
- Miguel et al. (2016) Eder Miguel, Mathias Lepoutre, and Bernd Bickel. 2016. Computational design of stable planar-rod structures. ACM Trans. Graph. 35, 4, Article 86 (jul 2016), 11 pages. https://doi.org/10.1145/2897824.2925978
- Mori and Igarashi (2007) Yuki Mori and Takeo Igarashi. 2007. Plushie: an interactive design system for plush toys. In ACM SIGGRAPH 2007 Papers (San Diego, California) (SIGGRAPH ’07). Association for Computing Machinery, New York, NY, USA, 45–es. https://doi.org/10.1145/1275808.1276433
- Mueller et al. (2014) Stefanie Mueller, Sangha Im, Serafima Gurevich, Alexander Teibrich, Lisa Pfisterer, François Guimbretière, and Patrick Baudisch. 2014. WirePrint: 3D printed previews for fast prototyping. In Proceedings of the 27th Annual ACM Symposium on User Interface Software and Technology (Honolulu, Hawaii, USA) (UIST ’14). Association for Computing Machinery, New York, NY, USA, 273–280. https://doi.org/10.1145/2642918.2647359
- Nuernberger et al. (2016) Benjamin Nuernberger, Eyal Ofek, Hrvoje Benko, and Andrew D. Wilson. 2016. SnapToReality: Aligning Augmented Reality to the Real World. In Proceedings of the 2016 CHI Conference on Human Factors in Computing Systems (San Jose, California, USA) (CHI ’16). Association for Computing Machinery, New York, NY, USA, 1233–1244. https://doi.org/10.1145/2858036.2858250
- Park et al. (2022) Keunwoo Park, Conrad Lempert, Muhammad Abdullah, Shohei Katakura, Jotaro Shigeyama, Thijs Roumen, and Patrick Baudisch. 2022. FoolProofJoint: Reducing Assembly Errors of Laser Cut 3D Models by Means of Custom Joint Patterns. In CHI Conference on Human Factors in Computing Systems. ACM, New Orleans LA USA, 1–12. https://doi.org/10.1145/3491102.3501919
- Pei et al. (2022) Siyou Pei, Alexander Chen, Jaewook Lee, and Yang Zhang. 2022. Hand interfaces: Using hands to imitate objects in ar/vr for expressive interactions. In Proceedings of the 2022 CHI conference on human factors in computing systems. 1–16.
- Roumen et al. (2019) Thijs Roumen, Jotaro Shigeyama, Julius Cosmo Romeo Rudolph, Felix Grzelka, and Patrick Baudisch. 2019. SpringFit: Joints and Mounts that Fabricate on Any Laser Cutter. In Proceedings of the 32nd Annual ACM Symposium on User Interface Software and Technology. ACM, New Orleans LA USA, 727–738. https://doi.org/10.1145/3332165.3347930
- Schumacher et al. (2015) Christian Schumacher, Bernd Bickel, Jan Rys, Steve Marschner, Chiara Daraio, and Markus Gross. 2015. Microstructures to control elasticity in 3D printing. ACM Trans. Graph. 34, 4, Article 136 (jul 2015), 13 pages. https://doi.org/10.1145/2766926
- Sethapakdi et al. (2021) Ticha Sethapakdi, Daniel Anderson, Adrian Reginald Chua Sy, and Stefanie Mueller. 2021. Fabricaide: Fabrication-Aware Design for 2D Cutting Machines. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (CHI ’21). ACM. https://doi.org/10.1145/3411764.3445345
- Skouras et al. (2014) Mélina Skouras, Bernhard Thomaszewski, Peter Kaufmann, Akash Garg, Bernd Bickel, Eitan Grinspun, and Markus Gross. 2014. Designing inflatable structures. ACM Trans. Graph. 33, 4, Article 63 (jul 2014), 10 pages. https://doi.org/10.1145/2601097.2601166
- Stava et al. (2012) Ondrej Stava, Juraj Vanek, Bedrich Benes, Nathan Carr, and Radomír Měch. 2012. Stress relief: improving structural strength of 3D printable objects. ACM Trans. Graph. 31, 4, Article 48 (jul 2012), 11 pages. https://doi.org/10.1145/2185520.2185544
- Stemasov et al. (2024) Evgeny Stemasov, Simon Demharter, Max Rädler, Jan Gugenheimer, and Enrico Rukzio. 2024. pARam: Leveraging Parametric Design in Extended Reality to Support the Personalization of Artifacts for Personal Fabrication. In Proceedings of the CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 337, 22 pages. https://doi.org/10.1145/3613904.3642083
- Stemasov et al. (2020) Evgeny Stemasov, Tobias Wagner, Jan Gugenheimer, and Enrico Rukzio. 2020. Mix&Match: Towards Omitting Modelling Through In-situ Remixing of Model Repository Artifacts in Mixed Reality. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’20). Association for Computing Machinery, New York, NY, USA, 1–12. https://doi.org/10.1145/3313831.3376839
- Tojo et al. (2024) Kenji Tojo, Ariel Shamir, Bernd Bickel, and Nobuyuki Umetani. 2024. Fabricable 3D Wire Art. In ACM SIGGRAPH 2024 Conference Proceedings (SIGGRAPH ’24). https://doi.org/10.1145/3641519.3657453
- Vanek et al. (2014) Juraj Vanek, JA Garcia Galicia, Bedrich Benes, R Měch, N Carr, Ondrej Stava, and GS Miller. 2014. Packmerger: A 3d print volume optimizer. In Computer Graphics Forum, Vol. 33. Wiley Online Library, 322–332.
- Wang et al. (2015) Weiming Wang, Haiyuan Chao, Jing Tong, Zhouwang Yang, Xin Tong, Hang Li, Xiuping Liu, and Ligang Liu. 2015. Saliency-preserving slicing optimization for effective 3D printing. In Computer Graphics Forum, Vol. 34. Wiley Online Library, 148–160.
- Weichel et al. (2014) Christian Weichel, Manfred Lau, David Kim, Nicolas Villar, and Hans W. Gellersen. 2014. MixFab: a mixed-reality environment for personal fabrication. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Toronto, Ontario, Canada) (CHI ’14). Association for Computing Machinery, New York, NY, USA, 3855–3864. https://doi.org/10.1145/2556288.2557090
- Xu et al. (2018) Hongyi Xu, Espen Knoop, Stelian Coros, and Moritz Bächer. 2018. Bend-it: design and fabrication of kinetic wire characters. ACM Trans. Graph. 37, 6, Article 239 (dec 2018), 15 pages. https://doi.org/10.1145/3272127.3275089
- Yang et al. (2021) Zhijin Yang, Pengfei Xu, Hongbo Fu, and Hui Huang. 2021. WireRoom: model-guided explorative design of abstract wire art. ACM Trans. Graph. 40, 4, Article 128 (jul 2021), 13 pages. https://doi.org/10.1145/3450626.3459796
- Yeh et al. (2022) Chih-Kuo Yeh, Thi-Ngoc-Hanh Le, Zhi-Ying Hou, and Tong-Yee Lee. 2022. Generating Virtual Wire Sculptural Art from 3D Models. ACM Trans. Multimedia Comput. Commun. Appl. 18, 2, Article 51 (feb 2022), 23 pages. https://doi.org/10.1145/3475798
- Zhou et al. (2016) Tian Zhou, Maria E. Cabrera, Juan P. Wachs, Thomas Low, and Chandru Sundaram. 2016. A comparative study for telerobotic surgery using free hand gestures. J. Hum.-Robot Interact. 5, 2 (sep 2016), 1–28. https://doi.org/10.5898/JHRI.5.2.Zhou